OpenStudy (anonymous):

I know this is simple but (6^3a)^3, Isn't it correct to rise the power of 3 to 9, so I end up with 6^9 and a^3

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

a^2 = a*a (a^2)^3 = (a*a)*(a*a)*(a*a) = a^6